B is 50% faster than A. If A starts at 9 A.M. and B starts at 10 A.M. A travels at a speed of 50 km/hr.
If A and B are 300 kms apart, The time when they meet when they travel in opposite direction is ?

Answer by mananth(16949)

A 50 km/h
...
50% of A = 25 km/h
...
B travels at 75km/h
..
B starts at 10 am.
A starts at 9 am
..
A travels 50 km in 1 hour
..
when B starts they are 300 -50 km apart.
=250km.
Speed together = 50+75 = 125 km/h
Time = distance /speed
Time =250/125
Time = 2hours to meet.
They meet at 12.00 noon